# Diffusion Bridge Implicit Models

### v2

@@ Abstract
Denoising diffusion bridge models (DDBMs) are a powerful variant of diffusion models for interpolating between two arbitrary paired distributions given as endpoints. Despite their promising performance in tasks like image translation, DDBMs require a computationally intensive sampling process that involves the simulation of a (stochastic) differential equation through hundreds of network evaluations. In this work, we take the first step in fast sampling of DDBMs without extra training, motivated by the well-established recipes in diffusion models. We generalize DDBMs via a class of non-Markovian diffusion bridges defined on the discretized timesteps concerning sampling, which share the same marginal distributions and training objectives, and give rise to generative processes ranging from stochastic to deterministic, resulting in diffusion bridge implicit models (DBIMs). DBIMs are not only up to 25$\times$ faster than the vanilla sampler of DDBMs but also induce a novel, simple, and insightful form of ordinary differential equation (ODE) which inspires high-order numerical solvers. Moreover, DBIMs maintain the generation diversity in a distinguished way, by using a booting noise in the initial sampling step, which enables faithful encoding, reconstruction, and semantic interpolation in image translation tasks. Code is available at https://github.com/thu-ml/DBIM.
@@ 1 Introduction
In this work, we present diffusion bridge implicit models (DBIMs) for accelerated sampling of DDBMs, motivated by the prominent denoising diffusion implicit models (DDIMs) (Song et al., 2021a). Firstly, we explore the continuous-time forward process of DDBMs on discretized timesteps which are exactly the sampling timesteps, revealing them to be Markovian diffusion bridges. We demonstrate that these bridges can be generalized to a series of non-Markovian processes controlled by a variance parameter, while maintaining identical marginal distributions and training objectives as DDBMs. Secondly, the induced reverse generative processes correspond to sampling procedures of varying levels of stochasticity, including deterministic ones. Consequently, DBIMs can be viewed as a bridge counterpart and extension of DDIMs. Furthermore, in the continuous time limit, DBIMs can induce a novel form of ordinary differential equation (ODE), which is linked to the probability flow ODE (PF-ODE) in DDBMs while being simpler and significantly more efficient. The induced ODE also facilitates novel high-order numerical methods for faster convergence.
@@ 3.1 Non-Markovian Diffusion Bridges as Forward Process
where $\rho_{n}$ is the $n$-th element of $\rho$ satisfying $\rho_{N-1}=c_{t_{N-1}}$. In the above expression, we reuse the notations of $a_{t},b_{t},c_{t}$ in Eqn. (6), which are known coefficients determined by the diffusion noise schedule. Intuitively, this decreases the variance (noise level) of the bridge (Eqn. (6)) while incorporating additional noise components from the last step. Under this construction, we can prove that $q^{(\rho)}$ maintains consistency in marginal distributions with $q$ defined in the forward process Eqn. (5).
@@
The definition of $q^{(\rho)}$ in Eqn. (10) represents the inference process, since it is factorized as the distribution of $\bm{x}_{t_{n}}$ given $\bm{x}_{t_{n+1}}$ at the previous timestep. Conversely, the forward process of $q^{(\rho)}(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{t_{n}},\bm{x}_{T})$ can be induced by Bayes’ rule (Appendix C.1). As $\bm{x}_{t_{n+1}}$ in $q^{(\rho)}$ can simultaneously depend on $\bm{x}_{t_{n}}$ and $\bm{x}_{0}$, we refer to it as non-Markovian diffusion bridges, in contrast to Markovian ones (such as Brownian bridges, and the diffusion bridge defined by the forward SDE in Eqn. (5)) which should satisfy $q(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{t_{n}},\bm{x}_{T})=q(\bm{x}_{t_{n+1}}|\bm{x}_{t_{n}},\bm{x}_{T})$.
@@ 3.2 Reverse Generative Process and Equivalent Training Objective
Eqn. (10) can be naturally transformed into a parameterized and learnable generative model, by replacing the unknown $\bm{x}_{0}$ in Eqn. (10) with a data predictor $\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$. Intuitively, $x_{t}$ on the diffusion bridge is a weighted mixture of $\bm{x}_{T},\bm{x}_{0}$ and some random Gaussian noise according to Eqn. (6), where the weightings $a_{t},b_{t},c_{t}$ are determined by the timestep $t$. The network $\bm{x}_{\theta}$ is trained to recover the clean data $\bm{x}_{0}$ given $\bm{x}_{t},\bm{x}_{T}$ and $t$.
@@ 4.2 Connection to Probability Flow ODE
We discuss some special cases of DBIM and their connection to related works in Appendix C.3.

### v3

@@ Abstract
Denoising diffusion bridge models (DDBMs) are a powerful variant of diffusion models for interpolating between two arbitrary paired distributions given as endpoints. Despite their promising performance in tasks like image translation, DDBMs require a computationally intensive sampling process that involves the simulation of a (stochastic) differential equation through hundreds of network evaluations. In this work, we take the first step in fast sampling of DDBMs without extra training, motivated by the well-established recipes in diffusion models. We generalize DDBMs via a class of non-Markovian diffusion bridges defined on the discretized timesteps concerning sampling, which share the same marginal distributions and training objectives, give rise to generative processes ranging from stochastic to deterministic, and result in diffusion bridge implicit models (DBIMs). DBIMs are not only up to 25$\times$ faster than the vanilla sampler of DDBMs but also induce a novel, simple, and insightful form of ordinary differential equation (ODE) which inspires high-order numerical solvers. Moreover, DBIMs maintain the generation diversity in a distinguished way, by using a booting noise in the initial sampling step, which enables faithful encoding, reconstruction, and semantic interpolation in image translation tasks. Code is available at https://github.com/thu-ml/DiffusionBridge.
@@ 1 Introduction
This work represents the first pioneering effort toward accelerated sampling of DDBMs. As suggested by well-established recipes in diffusion models, training-free accelerations of diffusion sampling primarily focus on reducing stochasticity (e.g. the prominent denoising diffusion implicit models, DDIMs) and utilizing higher-order information (e.g. high-order solvers). We present diffusion bridge implicit models (DBIMs) as an approach that explores both aspects within the diffusion bridge framework. Firstly, we investigate the continuous-time forward process of DDBMs on discretized timesteps and generalize them to a series of non-Markovian diffusion bridges controlled by a variance parameter, while maintaining identical marginal distributions and training objectives as DDBMs. Secondly, the induced reverse generative processes correspond to sampling procedures of varying levels of stochasticity, including deterministic ones. Consequently, DBIMs can be viewed as a bridge counterpart and extension of DDIMs. Furthermore, in the continuous time limit, DBIMs can induce a novel form of ordinary differential equation (ODE), which is linked to the probability flow ODE (PF-ODE) in DDBMs while being simpler and significantly more efficient. The induced ODE also facilitates novel high-order numerical diffusion bridge solvers for faster convergence.
@@ 3.1 Non-Markovian Diffusion Bridges as Forward Process
where $\rho_{n}$ is the $n$-th element of $\rho$ satisfying $\rho_{N-1}=c_{t_{N-1}}$, and $a_{t},b_{t},c_{t}$ are terms related to the noise schedule, as defined in the original diffusion bridge (Eqn. (6)). Intuitively, this decreases the variance (noise level) of the bridge while incorporating additional noise components from the last step. Under this construction, we can prove that $q^{(\rho)}$ maintains consistency in marginal distributions with the original forward process $q$ governed by Eqn. (5).
@@
The definition of $q^{(\rho)}$ in Eqn. (10) represents the inference process, since it is factorized as the distribution of $\bm{x}_{t_{n}}$ given $\bm{x}_{t_{n+1}}$ at the previous timestep. Conversely, the forward process $q^{(\rho)}(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{t_{n}},\bm{x}_{T})$ can be induced by Bayes’ rule (Appendix C.1). As $\bm{x}_{t_{n+1}}$ in $q^{(\rho)}$ can simultaneously depend on $\bm{x}_{t_{n}}$ and $\bm{x}_{0}$, we refer to it as non-Markovian diffusion bridges, in contrast to Markovian ones (such as Brownian bridges, and the diffusion bridge defined by the forward SDE in Eqn. (5)) which should satisfy $q(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{t_{n}},\bm{x}_{T})=q(\bm{x}_{t_{n+1}}|\bm{x}_{t_{n}},\bm{x}_{T})$.
@@ 3.2 Reverse Generative Process and Equivalent Training Objective
Eqn. (10) can be naturally transformed into a parameterized and learnable generative model, by replacing the unknown $\bm{x}_{0}$ in Eqn. (10) with a data predictor $\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$. Intuitively, $\bm{x}_{t}$ on the diffusion bridge is a weighted mixture of $\bm{x}_{T},\bm{x}_{0}$ and some random Gaussian noise according to Eqn. (6), where the weightings $a_{t},b_{t},c_{t}$ are determined by the timestep $t$. The network $\bm{x}_{\theta}$ is trained to recover the clean data $\bm{x}_{0}$ given $\bm{x}_{t},\bm{x}_{T}$ and $t$.
@@ 4.1 Diffusion Bridge Implicit Models
Related Work
We discuss related fast sampling and diffusion bridge works in Appendix A. We also discuss some special cases of DBIM and their connection to related works in Appendix C.3.